OpenStudy (anonymous):

One month Deandre rented 7 movies and 9 video games for a total of $78. The next month he rented 5 movies and 3 video games for a total of $36. Find the rental cost for each movie and each video game.

OpenStudy (anonymous):

Ok, you want to start by making two equations

OpenStudy (anonymous):

For our equations, x will be movies and y will be games

OpenStudy (anonymous):

First equation is 7x + 9y=78

OpenStudy (anonymous):

Have any idea what the second will be?

OpenStudy (anonymous):

@atlantaslast

OpenStudy (anonymous):

5x+3y=36

OpenStudy (anonymous):

Great! Now to solve we can use substitution, do you know how to use that?

OpenStudy (anonymous):

Not that well.

OpenStudy (anonymous):

That's fine, we'll work with the second one since it has smaller numbers.

OpenStudy (anonymous):

So the basic idea is we will turn one equation into a value, that will substitute it's value in the other equation

OpenStudy (anonymous):

that's a bunch of jibber jabber, but here is the math behind that statement

OpenStudy (anonymous):

y=12-5x

OpenStudy (anonymous):

Do you understand how I got to this from the original equation 5x+3y=36?

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

wait did you multiply by 4?

OpenStudy (anonymous):

Sorry I typed it wrong I meant to put y=12-5/3x

OpenStudy (anonymous):

No, first I subtracted 5x from both sides, and then divide both sides by 3

OpenStudy (anonymous):

the purpose is to isolate the y, have it on it's own

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

So now that we have y=12-5/3x, we can plug it into the other equation

OpenStudy (anonymous):

7x+9(12-5/3x)=78

OpenStudy (anonymous):

Do you understand what I just did?

OpenStudy (anonymous):

Yeah you plugged in the numbers to the equation

OpenStudy (anonymous):

I found that the value of y is 12-5/3x, so I substituted y for 12-5/3x in the other equation

OpenStudy (anonymous):

Now I want you to try to solve 7x+9(12-5/3x)=78 for x

OpenStudy (anonymous):

7x+106.3=78 right?

OpenStudy (anonymous):

No, we still don't know what x is.. so you would get 7x+108-15x=78

OpenStudy (anonymous):

which becomes 108-8x=78

OpenStudy (anonymous):

and then you subtract 8 from each side correct?

OpenStudy (anonymous):

i mean add

OpenStudy (anonymous):

not yet, first subtract 108 from each side, you need to isolate the x

OpenStudy (anonymous):

so you get -8x=-30

OpenStudy (anonymous):

now divide both sides by -8 and get x=15/4

OpenStudy (anonymous):

yep, when you actually divide 15 by 4 you get 3.75 so every movie costs $3.75 to rent

OpenStudy (anonymous):

now that we have that, we plug it into one of the originals, i'll use 5x+3y=36 since it's smaller numbers

OpenStudy (anonymous):

plug in 3.75 for x... 5(3.75)+3y=36

OpenStudy (anonymous):

and then you get 18.75+3y=36 right

OpenStudy (anonymous):

yep, now subtract 18.75 from both sides

OpenStudy (anonymous):

you should get 3y=17.25

OpenStudy (anonymous):

yep i got that

OpenStudy (anonymous):

divide both by 3 and get y=5.75

OpenStudy (anonymous):

so every movie is $3.75, every game is $5.75

OpenStudy (anonymous):

Thanks very much :)

OpenStudy (anonymous):

no problem!